OpenStudy (anonymous):

a. If the cost of printing the novel includes a base cost of $1,550 plus $4 per book, write a function that represents the cost of printing x copies of the book.

OpenStudy (anonymous):

1,550 + (4x)

OpenStudy (anonymous):

Or rather Y = 1,550 + (4x).

OpenStudy (anonymous):

can help a little more please

OpenStudy (anonymous):

b. If you want to cover the printing costs by printing and selling only 50 copies of the book, at what price must you sell the book?

OpenStudy (anonymous):

Well we have to work out the cost of printing 50 copies of the book. Stick '50' into the function as X and work it out.

OpenStudy (anonymous):

Y = 1,550 + (4x) 4 x 50 = 200 1550 + 200 is 1750.

OpenStudy (anonymous):

Then we have to divide £1750 by 50 to work out how much we need to sell for to cover the cost.

OpenStudy (anonymous):

Thirty-five quid.

OpenStudy (anonymous):

what

OpenStudy (anonymous):

can you do some more ill do new posts for medals if you lke

OpenStudy (anonymous):

Sure, I'll help some more.

OpenStudy (anonymous):

new post?